OpenStudy (anonymous):

If (x)=9-2|x+2|, find f (-7) (points: 1) A. -1 B. -15 C. 1 D. 19 Help? Please? :-)

jimthompson5910 (jim_thompson5910):

f(x)=9-2|x+2| f(-7)=9-2|-7+2| ... replace every copy of x with -7, then evaluate f(-7) = ???